OpenStudy (i_ask_alot):

The area of a circle is 615.44 in^2. What is the diameter of the circle? Use 3.14 for π.

OpenStudy (i_ask_alot):

@Michele_Laino

OpenStudy (michele_laino):

we can apply this formula: \[\huge A = \frac{{\pi {d^2}}}{4}\] where \(A\) is the area, and \(d\) is the diameter

OpenStudy (michele_laino):

If we write the inverse formula, we get: \[\huge d = \sqrt {\frac{{4A}}{\pi }} \]

OpenStudy (i_ask_alot):

is the answer 14 in

OpenStudy (michele_laino):

or: \[\huge d = \sqrt {\frac{{4A}}{\pi }} = \sqrt {\frac{{4A}}{{3.14}}} \]

OpenStudy (michele_laino):

I got a different value

OpenStudy (i_ask_alot):

mmm let me see

OpenStudy (i_ask_alot):

20 in

OpenStudy (michele_laino):

hint: here is the next step: \[\huge \begin{gathered} d = \sqrt {\frac{{4A}}{\pi }} = \sqrt {\frac{{4A}}{{3.14}}} = \hfill \\ \hfill \\ = \sqrt {\frac{{4 \times 615.44}}{{3.14}}} = \sqrt {\frac{{2461.76}}{{3.14}}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (i_ask_alot):

ok is it 28

OpenStudy (michele_laino):

that's right!

OpenStudy (i_ask_alot):

thanks

OpenStudy (michele_laino):

:)